Suppose that 69% of all college seniors have a job prior to graduation. If a random sample of 50 college seniors is taken, approximate the probability that more than 37 have a job prior to graduation.
Use the normal approximation to the binomial with a correction for continuity.

Answer by Boreal(15235)

mean is np=50*0.69=34.5
variance is np(1-p)=34.5*0.31=10.695
sd is sqrt(V)=3.27
want a z-value. More than 37 means start greater than 37 and the continuity correction factor has that at 37.5 and greater
z>(37.5-34.5)/3.27
z>3/3.27 =0.92
probability z>0.92 is 0.1795
(exact value is 0.1803)